# Long-Tailed Continual Learning For Visual Food Recognition

### v1

@@ V Experiment
In this section, we evaluate our proposed long-tailed continual learning framework for visual food recognition as illustrated in Section IV. Specifically, we first introduce the experimental setup including the split of datasets and implementation detail described in Section V-A and V-B, respectively. Then we compare our method with existing work in Section V-C and conduct an ablation study to show the effectiveness of each individual component in Section V-D. Finally, we discuss potential techniques that can boost the performance for real world food related applications in Section V-E.
@@ V-A Datasets
We conduct experiments on four benchmark long-tailed food image datasets including: (1) Food101-LT [14], (2) VFN-LT [14], (3) VFN-INSULIN and (4) VFN-T2D where the latter two datasets are introduced in Section III.
Food101-LT is the long-tailed version of Food-101 [12] constructed by applying Pareto distribution [71] with the power ratio $\alpha=6$. We randomly partition the 101 food classes into 5, 10, and 20 tasks for continual learning. Therefore, each task comprises 20, 10, and 5 new classes, respectively, with the exception of the first task, which includes one additional class. The test set is kept as balanced with 125 images per class.
VFN-LT, VFN-INSULIN and VFN-T2D are all the long-tailed version of VFN [9] but constructed based on the food consumption frequency of different populations including (i) healthy people (ii) insulin takers and (iii) individuals with type 2 diabetes without taking insulin, respectively. We randomly divide the 74 food classes into $N=7$ tasks with the first task contain 14 new classes and the remaining tasks contain 10 new classes for continual learning. The test set is balanced with 25 images per class. To facilitate an equatable analysis, we use the same testing data in VFN-LT, VFN-INSULIN and VFN-T2D, which is balanced as 25 samples per class with total 1,850 images.
@@ V-B Implementation Detail
Our implementation of neural networks are based on Pytorch framework [72] and we apply the ResNet-18 from scratch as the backbone network for experiments for all datasets. The ResNet implementation follows the setting suggested in [11]. We train 90 epochs for each new task with the learning rate starts from 0.1 and decreased with ratio $1/10$ for every 30 epochs. The batch size is set as 128 and we apply the stochastic gradient descent (SGD) optimizer with weight decay $0.0001$. The memory budget is set as $\mathcal{M}=20$ to store at most 20 images per food class in the memory buffer.
Evaluation protocol: We use Top-1 classification accuracy as the evaluation metric and evaluate the updated model after learning each new task $\mathcal{T}^{i}$ on test data belonging to all classes seen so far $C_{1:i}$. Besides, we report the last step accuracy $A_{L}$ defined as the classification accuracy on the entire test set after learning all $N$ tasks and the average accuracy $A_{M}$ calculated by averaging the accuracy obtained after learning each new task, which shows the overall performance for the entire continual learning procedure. Each experiment is run five times and the average performance is presented.
Existing methods: We compare our proposed framework with existing continual learning methods including LwF [55], EWC [54], iCaRL [60], EEIL [61], LwM [70], IL2M [19], BiC [62], LUCIR [57] and PODNet [58]. Also, we apply the 2-stage framework for long-tailed continual learning in [17] and incorporate it into three existing methods including EEIL-2stage, LUCIR-2stage and PODNet-2stage for comparisons.
Figure 6: Results on VFN-LT, VFN-INSULIN, VFN-T2D and Food101-LT with different number of tasks $N$. Each marker represents the Top-1 classification accuracy evaluated on all classes seen so far after learning each task.
@@ V-C Comparisons With Existing Methods
Table II summarizes the results on Food101-LT, VFN-LT, VFN-INSULIN and VFN-T2D in terms of the last step accuracy $A_{L}$ and the average accuracy $A_{M}$. We observe noticeable improvements for both $A_{L}$ and $A_{M}$ of our proposed method by comparing with existing work especially on VFN-LT, VFN-INSULIN and VFN-T2D, which contains a heavier tail with a larger imbalance ratio as introduced in Section III. For example, we achieve around $5\%$ increase of average accuracy $A_{M}$ and $2\%$ increase of the last step accuracy $A_{L}$ on VFN-INSULIN compared with the 2-stage framework even without requiring the decoupled training process. Furthermore, we obtain the best performance on Food101-LT with different number of tasks $N\in\{5,10,20\}$. We notice both $A_{L}$ and $A_{M}$ accuracy will drop as the total number of tasks $N$ increase since we need to address the catastrophic forgetting to maintain the learned knowledge at each learning phase of new tasks except for the fist task.
Figure 6 shows the results of top-1 classification accuracy evaluated on all the classes seen so far after learning each new task. Our method achieve promising performance at each learning phase of new task. In addition, different from conventional continual learning where the accuracy usually drops monotonously overtime when new tasks arrive due to the forgetting of previous knowledge, we notice that the performance may even increase after learning a new task in the long-tailed scenario. For example on Food101-LT with $N=5$, we observe the increase of accuracy for most methods after learning the third task. This occurs because the number of training samples varies significantly among different tasks in long-tailed continual learning where the model gains better knowledge for tasks with a larger number of training images. However, this imposes new challenges of learning from class-imbalanced data for each new task and also hyper-parameter tuning (e.g. the factor of knowledge distillation term as in Equation 4). In this work, we not only address the catastrophic forgetting by introducing a feature-based knowledge distillation and an effective exemplar augmentation technique but also alleviate the class-imbalance by integrating the balanced softmax with an adaptive ratio to adjust the impact of each loss term as illustrated in Equation 11 and strike a balance of stability-plasticity.
TABLE III: Ablation study on Food101-LT, VFN-LT, VFN-INSULIN and VFN-T2D in terms of average accuracy $A_{M}$.
Food101-LT VFN-LT VFN-INSULIN VFN-T2D
@@
5.90 8.79 10.55 12.21 11.47 11.93
17.42 15.83 13.96 22.53 21.19 20.73
13.27 12.99 11.64 16.73 15.41 14.93
16.52 14.20 12.02 22.96 21.33 21.05
19.31 17.26 15.49 24.18 22.31 22.19
21.83 19.25 17.43 29.33 27.51 28.17
@@ V-D Ablation Study
We also evaluate our proposed CAM-CutMix by replacing it with existing data augmentation based methods including the CutMix [22] based approaches: (a) the original CutMix used in CMO [21], (b) Visual-Multi CutMix (VM-CMO[14], (c) SnapMix [50] and the Mixup [73] based approach: (d) D-Mixup [41]. We conduct experiments on VFN-LT and Food101-LT with $N=10$ as shown in Table IV. Generally, the CutMix based approaches work better in long-tailed continual learning scenarios than D-Mixup, which is usually applied in multi-label recognition scenarios. In addition, the SnapMix achieve a slightly better performance than CMO and VM-CMO as it also considers the class-activation map (CAM) when generating mixed labels. Our method achieves the best performance as it not only preserves the most important regions based on CAM but also enables seamless CutMix, rather than relying on a randomly generated bounding box. The example augmented food images using VFN-LT are shown in Figure 7. Note that we do not visualize SnapMix [50] as it has the same synthetic image as in CMO [21] but with a different mixed label.
@@ V-E Discussions
Despite the performance improvements our proposed framework demonstrates in comparison to existing methods as shown in Table II, the deployment in real-world applications still remains challenging based on the current classification accuracy. Therefore, in this section we discuss potential techniques that could be applied to boost the performance including (1) increaseing the memory buffer capacity to store more exemplar images for knowledge replay and (2) performing transfer learning by pre-training the backbone on large scale image datasets.
V-E1 Memory buffer capacity
@@
As one of the most efficient techniques to address catastrophic forgetting, the performance of knowledge replay greatly relies on the capacity of memory buffer (i.e. and how many exemplar images can be stored). In this part, we evaluate the long-tailed continual learning performance by varying the memory buffer capacity $\mathcal{M}\in\{10,20,30,40,50,100\}$. Figure 8 shows the results in terms of average accuracy $A_{M}$ on Food101-LT ($N=10$) and VFN-LT where we observe consistent performance improvements by increasing the memory capacity. However, the memory buffer capacity is a significant constraint for continual learning in real world application as it requires larger memory storage and also poses challenges related to privacy concerns when storing original images as exemplars. Additionally, we observe a performance bottleneck where increasing memory capacity does not substantially boost performance. This is predominantly due to dual challenges of catastrophic forgetting and class-imbalance problems that arise in the long-tailed continual learning scenario.
@@
Figure 8: Average accuracy ($A_{M}$) on Food101-LT ($N=10$) and VFN-LT by varying the memory buffer capacity $\mathcal{M}\in\{10,20,30,40,50,100\}$.
V-E2 Transfer learning
@@
Applying the deep models pre-trained on large scale image datasets as backbone a network has been a common strategy to enhance performance across a multitude of vision tasks [74, 75, 13]. In this part, we evaluate the efficacy by leveraging a variety of pre-trained models within the context of long-tailed continual learning for visual food recognition. Formally, we consider different network structures with various depth such as ResNet-50 [11], MobileNet [76, 77], EfficientNet [78, 79] Vision Transformers (ViT) [80, 81, 82] and its variants including DeiT [81] and Swin [82] transformers. In addition, we leverage both ImageNet-1K [83] and ImageNet-21K [84] as the pre-training datasets. ImageNet-1K contains 1,000 classes of general objects, which is the subset of full ImageNet-21K that contains 21,841 classes with over 14,197,122 training images. The results on VFN-LT in terms of average accuracy $A_{M}$ is shown in Figure 9. We observe significant performance improvements for over $20\%$ by using pre-trained models on large-scale datasets compared to our results in Table II with a model from scratch. This is mainly attributed to the fact that pre-training enhances the feature extraction capabilities of the backbone network, thereby enabling it to procure the most discriminative features that are essential for downstream tasks. In addition, pre-training on larger scale datasets that contain more images and classes is able to achieve higher accuracy. However, there is trade-off between the computation complexity and the performance where the increase of model parameters would require longer training time and higher computation capability, which may not be practical for specific real world applications with limited resources available. Note that we intentionally refrain from utilizing food datasets for pre-training in this part to prevent potential overlap with any food class in VFN [9]. However, it’s reasonable to anticipate a more substantial performance enhancement if the pre-training were conducted on large-scale food datasets such as Food2K [13].
Figure 9: Average accuracy ($A_{M}$) on VFN-LT by leveraging pre-trained models.

### v2

@@ V Experiment
In this section, we evaluate our proposed long-tailed continual learning framework for visual food recognition as illustrated in Section IV. Specifically, we first introduce the experimental setup including the split of datasets and implementation detail described in Section V-A and V-B. Then we compare our method with existing work in Section V-C and conduct an ablation study to show the effectiveness of each individual component in Section V-D. Finally, we discuss potential techniques that can boost the performance of real-world food-related applications in Section V-E.
@@ V-A Datasets
Food101-LT is the long-tailed version of Food-101 [10], created using the Pareto distribution [81] with the power ratio of $\alpha=6$. We randomly partition the 101 food classes into 5, 10, and 20 tasks for continual learning, where each task introduces 20, 10, and 5 new classes, respectively, except the first task with one extra class. The test set is kept as balanced with 125 images per class.
VFN-LT is a long-tailed version of VFN [7] based on food consumption frequency of healthy people. The 74 food classes are split into 7 tasks, with the first task containing 14 new classes and the remaining tasks containing 10 new classes. The test set has 25 images per class.
VFN186-LT, VFN186-INSULIN and VFN186-T2D are long-tailed versions of VFN186. VFN186-LT is created similarly to VFN-LT, while VFN186-INSULIN and VFN186-T2D are based on food consumption frequencies of insulin takers and individuals with type 2 diabetes without insulin. We divide 186 food classes into $N=9$ tasks with the first task containing 26 new classes and the rest containing 20. To facilitate an equatable analysis, we use the same testing data in VFN186-LT, VFN186-INSULIN and VFN186-T2D, which is balanced with 25 samples per class, totaling 4,650 images.
ImageNetSubset-LT is a subset of ImageNet [82]. We follow [22, 80] to select 100 classes and apply the long-tailed transformation. Specifically, we randomly remove training samples following an imbalance factor of $\rho=n_{\text{max}}/n_{\text{min}}=100$, where $n_{\text{max}}$ and $n_{\text{min}}$ denote the maximum and minimum number of training samples per class, respectively. The dataset is split evenly into $N=10$ task and $N=20$ tasks, where each task contains 10 and 5 new classes, respectively. The test set remains unchanged, preserving the original class-balanced distribution.
@@ V-B Implementation Detail
Our implementation of neural networks are based on the Pytorch framework and we apply the ResNet-18 from scratch as the backbone for all experiments. The ResNet implementation follows the setting suggested in [9]. We train each new task for 90 epochs with the learning rate starting from 0.1 and decreasing with a ratio of $1/10$ for every 30 epochs. The batch size is set to 128 and we apply the stochastic gradient descent (SGD) optimizer with a weight decay of $0.0001$. To ensure a fair comparison between our method and existing methods, we set the random seed to 1993, following [64, 83].
Exemplar Selection Strategy: To construct the memory buffer, we follow the benchmark setting in [80] and set the memory budget to $\mathcal{M}=20$, allowing the storage of at most 20 exemplars per class. We adopt herding algorithm [63] for exemplar selection, which selects samples closest to the class mean in the feature space as exemplars. For tailed classes are with fewer than $\mathcal{M}$ available samples, we employ our CAM-based exemplar augmentation strategy to generate synthetic exemplars, ensuring balanced memory utilization across all classes.
Evaluation protocol: We use Top-1 classification accuracy to evaluate the model after each task $\mathcal{T}^{i}$ on test data covering previously seen classes $C_{1:i}$. Besides, we report the average accuracy $A_{M}$, calculated by averaging the accuracy after each task, which shows the overall performance across the continual learning procedure. Each experiment is run five times and the average performance is presented.
Figure 5: Results on Food101-LT, VFN-LT, VFN186-LT, VFN186-INSULIN and VFN186-T2D with different number of tasks $N$. Each marker represents the Top-1 classification accuracy evaluated on all classes seen so far after learning each task.
@@ V-C Comparisons With Existing Methods
V-C1 Performance across Datasets
Table I summarizes the average accuracy $A_{M}$ on Food101-LT, VFN-LT, VFN186-LT, VFN186-INSULIN, VFN186-T2D, and ImageNetSubset-LT. Our method shows significant improvements, particularly on Food101, with different numbers of tasks $N\in\{5,10,20\}$, achieving approximately a 5% increase in accuracy. Moreover, there are enhancements on VFN186-LT, VFN186-INSULIN, and VFN186-T2D, which feature more imbalanced distributions as discussed in Section III. For example, on three long-tailed VFN186, we achieve about a $7\%$ increase over the 2-stage framework even without requiring a decoupled training process and a $3\%$ improvement compared to DGR. However, tends often decreases as the total number of tasks $N$ increases. Therefore, we need to address the catastrophic forgetting to maintain the learned knowledge at each learning phase of new tasks after the first task. However, improvements are not evident on VFN. Considering the imbalanced data and the differences between food images are much smaller than those in other scenarios, this type of tasks is more difficult, making it hard to achieve significant improvements in classification results. Additionally, we do not present average performance in general scenarios since the random seed is fixed. In this case, we only evaluate the algorithm itself, rather than its performance across various situations. To demonstrate the broader applicability of our method beyond food data, we conduct additional experiments on ImageNetSubset for long-tailed learning across diverse object categories. Despite being primarily designed for food-related tasks, our approach shows strong generalization capability, achieving competitive performance against state-of-the-art methods. These results underscore the robustness of our method to effectively handle class imbalance, suggesting the potential for visual recognition tasks across multiple domains.
V-C2 Visualization and Upper Bound
Figure 5 shows top-1 classification accuracy across all seen classes after each task and upper bounds for continual learning. We use the joint training results as the upper bounds. Specifically, we train models on the full long-tailed datasets using three different strategies: vanilla cross-entropy (CE), balanced softmax (BS)[84], and Context-rich Minority Oversampling (CMO)[85]. To ensure a fair comparison with the long-tailed continual learning experiments, the backbone and configurations are keep the same of other experiments. We report the final performance on the last task, as joint training is equivalent to training on the entire dataset at once. As shown in Figure 5, these three settings achieve the best overall performance on Food101-LT and VFN-LT. Notably, on VFN186, CE and CE+CMO even perform worse than continual learning methods. This is due to the severe class imbalance and the limited number of samples in certain categories. Notably, as discussed in [86], the sample counts of rare classes in VFN186 are significantly smaller than those in Food101-LT and VFN-LT, which exacerbates the performance drop on VFN186.
Our method achieves promising performance at each stage of the new task. Interestingly, in long-tailed scenarios, unlike conventional continual learning where accuracy typically declines over time, we observe improvements after learning new tasks sometimes. For example, on VFE186-LT, accuracy increases for LwF, BiC, LUCIR, PODNet, and EEIL-2Stage after the task $N=4$. This occurs because the number of training samples varies significantly among different tasks in long-tailed continual learning where the model gains better knowledge for tasks with a larger number of training images. Therefore, it is common for accuracy to improve when learning classes with larger sample sizes in long-tailed continual learning. However, it also imposes new challenges in handling class-imbalance across different tasks and hyper-parameter tuning (e.g. the knowledge distillation factor in Equation 3).
V-C3 Comparisons With Prompt-Based Continual Learning
We further compare our method with recent prompt-based continual learning approaches including L2P [68], DualPrompt [69], and CODA-Prompt [70]. We adopt ViT-B/16 [87] pretrained on ImageNet-21K [88] as our backbone architecture across all experiments. We train each new task for $20$ epochs with batch size of $48$ using SGD optimizer. The initial learning rate is set to $0.05$ and decays following a cosine scheduler. To ensure fair comparisons, we follow [68] to select the same amount of exemplars as used in our method. Each experiment is conducted five times, and the reported results represent the mean accuracy with standard deviation on Food101-LT for task numbers $N=10$ and $N=20$. As shown in Table II, our method consistently outperforms all prompt-based methods on Food101-LT for all tasks, and the relatively low standard deviation in our experimental results indicates that our method provides stable performance. These results indicate our strong adaptation ability to work with pre-trained backbones.
@@ V-C3 Comparisons With Prompt-Based Continual Learning
Food101-LT
$N=10$ $N=20$
L2P-R [68] $82.01\pm 0.89$ $75.58\pm 0.52$
DualPrompt [69] $83.33\pm 0.25$ $74.29\pm 1.41$
CODA-Prompt [70] $85.14\pm 0.39$ $78.72\pm 0.41$
Ours $86.52\pm 0.18$ $80.45\pm 0.48$
TABLE III: Ablation study on Food101-LT and VFN-LT in terms of average accuracy $A_{M}$.
  Food101-LT VFN-LT
[…]
5.90 8.79 10.55 12.21
17.42 15.83 13.96 22.53
13.27 12.99 11.64 16.73
16.52 14.20 12.02 22.96
19.31 17.26 15.49 24.18
21.83 19.25 17.43 29.33
@@ V-D1 Effectiveness of Core Components
V-D2 Effectiveness of CAM-CutMix
We further evaluate our proposed CAM-CutMix by replacing it with existing data augmentation based methods including the CutMix [28] based approaches: (a) the original CutMix used in CMO [27], (b) Visual-Multi CutMix (VM-CMO[12], (c) SnapMix [54] and the Mixup [89] based approach: (d) D-Mixup [47]. We conduct experiments on VFN-LT and Food101-LT with $N=10$ as shown in Table IV. Generally, the CutMix-based approaches work better in long-tailed continual learning scenarios than D-Mixup, which is usually applied in multi-label recognition scenarios. In addition, the SnapMix achieves a slightly better performance than CMO and VM-CMO as it also considers the class-activation map (CAM) when generating mixed labels. Our method achieves the best performance as it not only preserves the most important regions based on CAM but also enables seamless CutMix, rather than relying on a randomly generated bounding box. The example augmented food images using VFN-LT are shown in Figure 6. Note that we do not visualize SnapMix [54] as it has the same synthetic image as in CMO [27] but with a different mixed label.
@@ V-D2 Effectiveness of CAM-CutMix
V-D3 Robustness to Design Variations
To evaluate the robustness of our method, we conducte additional experiments by replacing key components with alternative approaches: (i) we substitute our herding exemplar selection with random selection (denoted as Random) for memory replay, (ii) we replace the cosine embedding loss with Mean Squared Error (MSE) loss in the knowledge distillation module, and (iii) we use Grad-CAM instead of the original CAM approach for generating attention maps. The results on Food101-LT with different task numbers ($N=5,10,20$) are shown in Table V, where the up arrow $\uparrow$ indicates an improvement over the original setting while the down arrow $\downarrow$ denotes a performance drop. The variations in performance across different configurations highlight the flexibility of our method, as it maintains stable accuracy regardless of the specific component used. Grad-CAM augmentation yields the most consistent improvements (e.g., $\uparrow 2.27$ at $N=10$), demonstrating its effectiveness in selecting semantic important regions. MSE loss results in mixed performance, with minor improvements at $N=10$ but drops at $N=5$ and $N=20$. Overall, these results confirm that our framework is adaptable and can accommodate different design choices while maintaining strong continual learning performance.
TABLE V: Ablation study on Food101-LT with different variants. Results are reported as average accuracy (%) across different task settings ($N=5,10,20$).
Food101-LT
$N=5$ $N=10$ $N=20$
Ours w/ Random 27.74$(\uparrow 0.22)$ 24.88$(\downarrow 0.24)$ 21.53$(\downarrow 0.19)$
Ours w/ MSE Loss 26.75$(\downarrow 0.77)$ 25.38$(\uparrow 0.26)$ 21.25$(\downarrow 0.47)$
Ours w/ Grad-CAM 28.83$(\uparrow 1.31)$ 27.39$(\uparrow 2.27)$ 23.08$(\uparrow 1.36)$
Ours (original) 27.52 25.12 21.72
@@ V-E Discussions
Despite the performance improvements our framework demonstrates compared to existing methods as shown in Table I, the deployment in real-world applications remains challenging due to current classification accuracy and computational complexity. Therefore, in this section, we (1) analyze the running time of different methods to assess computational efficiency, and discuss potential techniques that could be applied to boost the performance including (2) increasing the memory buffer capacity to store more exemplar images for knowledge replay and (3) performing transfer learning on our methods to evaluate its scalability across different pretraining.
V-E1 Computational Complexity
Regarding computational complexity, we record the time required for each model to be trained from scratch, including the time needed for testing. Our method takes 69 minutes to finish the whole training process. As seen in Figure 7, there is a 15-minute reduction compared to the recently released DGR [80] method, which requires 84 minutes, but it achieves superior classification performance, making it both time-efficient and effective. iCaRL [64] stands out as the fastest but at the cost of lower classification accuracy, particularly in handling datasets of long-tailed distribution. Three two-stage methods show robust classification accuracy but at the expense of significantly higher training times, which are all over three times to our approach’s. Our method strikes an optimal balance between computational efficiency and classification accuracy, outperforming all other methods when considering both aspects, which makes it particularly suitable for real-world applications.
Figure 7: Running time (min) comparison on VFN186-LT for different models.
V-E2 Memory buffer capacity
@@
As one of the most efficient techniques to address catastrophic forgetting, the performance of knowledge replay greatly relies on the capacity of the memory buffer (i.e. how many exemplar images can be stored). In this part, we evaluate the long-tailed continual learning performance by varying the memory buffer capacity $\mathcal{M}\in\{10,20,30,40,50,100\}$. To ensure fair class-wise representation under the long-tailed setting, we adopt a fixed number of exemplars per class. Table VI reports the average accuracy $A_{M}$ on Food101-LT ($N=10$) and VFN-LT. We observe consistent performance improvements as $\mathcal{M}$ increases. However, the memory buffer capacity is a significant constraint for continual learning in real-world applications as it requires larger memory storage and also poses challenges related to privacy concerns when storing original images as exemplars. Additionally, the gain saturates at different points depending on the dataset. For instance, Food101-LT continues to improve up to $\mathcal{M}=40$, while VFN-LT shows marginal improvement beyond $\mathcal{M}=20$. This suggests a dataset-dependent trade-off between buffer size and performance. Moreover, the performance bottleneck is predominantly due to dual challenges of catastrophic forgetting and class-imbalance problems that arise in the long-tailed continual learning scenario.
@@
TABLE VI: Average accuracy ($A_{M}$) on Food101-LT ($N=10$) and VFN-LT by varying the memory buffer capacity $\mathcal{M}\in\{10,20,30,40,50,100\}$.
Buffer Size 10 20 30 40 50 100
Food101-LT 16.96 19.69 20.71 22.15 23.14 24.30
VFN-LT 25.71 29.33 30.84 31.25 31.89 32.55
V-E3 Variants of Backbones and Pre-training Datasets
@@
Applying the deep models pre-trained on large-scale image datasets as the backbone is a common strategy to enhance performance in many vision tasks [90, 11]. In this part, we investigate how our method performs when applied to different backbone architectures and pre-training datasets. Instead of modifying the learning paradigm, we analyze whether our approach remains effective across various network structures and different levels of pre-training. We consider backbones with various depth including ResNet-50 [9], MobileNet [91], EfficientNet [92] Vision Transformers (ViT) [87] and its variants DeiT [93] and Swin [94] transformers. In addition, we leverage ImageNet-1K [95] and ImageNet-21K [88] as the pre-training datasets. ImageNet-1K contains 1,000 classes of general objects, which is the subset of full ImageNet-21K that contains 21,841 classes with over 14,197,122 training images. The VFN-LT results in average accuracy $A_{M}$ are shown in Table VII. We observe over $20\%$ performance improvements by using pre-trained models on large-scale datasets compared to our results in Table I with a model from scratch. It manifests that pre-training enhances the backbone network’s feature extraction capabilities, thereby yielding the most discriminative features essential for downstream tasks. In addition, pre-training on larger-scale datasets with more images and classes makes higher accuracy. However, there is a trade-off between the computation complexity and the performance where the increase of model parameters would require longer training time and higher computation capability, which may not be practical for specific real-world applications with limited resources. Note that we intentionally refrain from utilizing food datasets for pre-training in this part to prevent potential overlap with any food class in VFN [7], though there may be a more substantial performance enhancement if pre-trained on large-scale food datasets such as Food2K [11].
TABLE VII: Average accuracy ($A_{M}$) on VFN-LT by leveraging pre-trained models.
Model MobileNet ResNet EfficientNet ViT DeiT Swin
Parameters (10M) 1.8 2.5 5.4 8.2 8.5 8.8
ImageNet-1K 54.99 56.73 61.78 59.75 61.41 63.17
ImageNet-21K 56.64 58.92 63.83 64.79 65.01 71.18